The Centers for Disease Control and Prevention estimate that approximately 100,000 Americans suffer nonfatal dog- and cat-assa) Suppose one injured individual is selected at random. What is the probability that this person is 75-or-over and
experienced a dog-related injury?


b) Suppose one injured individual is selected at random. What is the probability that this person is between 55 and
74 years old?


c) Suppose one injured individual is selected at random. What is the probability that this individual is at least 55
years old (i.e., is 55 years old or older)?


d) Suppose one injured person is selected at random. What is the probability that he or she is 75-or-over with a
cat-related injury OR 14-or-under with a dog-related injury?


e) Suppose one injured individual is selected at random. What is the probability that he or she is between 15 and
54 years old?


f) Suppose one individual is selected at random. What is the probability that he or she experienced a cat-related
injury?
Answer #1

Let p1 be the probability that a man or lady in age group 35-54 injured by a dog

p2 be the probability that a man or lady in age group 15-34 injured by a cat

then 0.15+0.16+p1+0.19+0.12+0.01+p2+0.04+0.02+0.02=1

hence p1+p2=1-0.71=0.29

p1=0.29-p2.hence the value of p1 and p2 lies in the interval[0,0.29

a) 0.12

b)let A be the event that injured person is between 55 and 74 years old.B1 be the event of injury by dog and B2 be the event of injury by cat

then P(A)= P(A/B1)*P(B1)+P(A/B2)*P(B2)

=0.19*(0.62+p1) +0.02*(0.09+p2)

=0.1196+0.19p1+0.02p2

hence P(A) lies in the interval [ 0.1254, 0.1747 ) as value of p1 and p2 lies in the interval[0,0.29

c) Atleast 55 include the age group 55-74 and 75 or over and let this event be B

p(Atleast 55)= P(B)=P(B/B1)*P(B1)+P(B/B2)*P(B2)

=0.31*(0.62+p1) +0.04*(0.09+p2)

=0.1958+0.31p1+0.04p2.

hence P(A) lies in the interval[0.199396.0.28571 as value of p1 and p2 lies in the interval[0,0.29

d) 0.02+0.15=0.17